OpenStudy (anonymous):

I am new to this environment and I need to know how to ask questions using the latex format. Could someone advise me please?

OpenStudy (anonymous):

In order to use latex, use \ [ to start and \ ] to end When using \ [ and \ ] make sure they are close together with no space. So to use latex you would do some thing like below \ [ \frac{2}{6} \ ] which would look like \[ \frac{2}{6} \] Remember not to have any space betwee the characters \ [ or \ ]

OpenStudy (anonymous):

\[ \frac{5}{45} \] wow is works.. lol

OpenStudy (anonymous):

thanks..

OpenStudy (anonymous):

Anytime and you are welcome